OpenStudy (anonymous):

Checking my answer for Let f(x) = 7x - 6 and g(x) = 3x + 4. Find f(x) + g(x).? would it be just 10x-2 or is it possible to simplify more?

Parth (parthkohli):

I guess there's nothing more

OpenStudy (anonymous):

Is it possible to simplify to 1/5

OpenStudy (asnaseer):

you could factor out a 2 as 2 divides into both 10x and 2

OpenStudy (inkyvoyd):

What do you mean, @Lovely95 ? Is that the answer given to you?

Parth (parthkohli):

Ahh yes we have the factorisation if needed.

OpenStudy (unklerhaukus):

\(f(x) = 7x - 6\) and \(g(x) = 3x + 4\). \[f(x) + g(x)=7x-6+3x+4\]\[\quad\qquad\qquad=10x-2\]\[\qquad\qquad\quad=2(5x-1)\] i dont think you can go any further

OpenStudy (anonymous):

ok...thanks!

OpenStudy (anonymous):

I shouldnt be solving for x?

OpenStudy (unklerhaukus):

is that what the question is asking? if so, go from here there is only two steps f(x)+g(x)=10x−2

OpenStudy (anonymous):

...1/5?

Parth (parthkohli):

This is a function, so you should really check if we have an equation or not. Here, no because it just tells to simplify.

OpenStudy (anonymous):

ok, im assuming you have to solve for x but it really doesnt say, thanks... :D

OpenStudy (unklerhaukus):

if the question is exactly what you wrote i dont think you have to solve for x

OpenStudy (inkyvoyd):

Lovely, you can't solve for x, because the equation is given in two variables (x and f(x) ). You would need a value of f(x) to solve for x.

OpenStudy (unklerhaukus):

\[x=\frac{f(x)+g(x)+2}{10}\]